# Accelerating Natural Gradient Descent for PINNs with Randomized Numerical Linear Algebra

### v2

@@ Abstract
Natural Gradient Descent (NGD) has emerged as a promising optimization algorithm for training neural network-based solvers for partial differential equations (PDEs), such as Physics-Informed Neural Networks (PINNs). However, its practical use is often limited by the high computational cost of solving linear systems involving the Gramian matrix. While matrix-free NGD methods based on the conjugate gradient (CG) method avoid explicit matrix inversion, the ill-conditioning of the Gramian significantly slows the convergence of CG. In this work, we extend matrix-free NGD to broader classes of problems than previously considered and propose the use of Randomized Nyström preconditioning to accelerate convergence of the inner CG solver. The resulting algorithm demonstrates substantial performance improvements over existing NGD-based methods on a range of PDE problems discretized using neural networks.
@@ 1 Introduction
Neural network-based approaches for solving partial differential equations (PDEs) have gained significant attention as promising alternatives to traditional numerical methods for both forward and inverse problems. A common strategy is to embed the governing PDE directly into the loss function, as in Physics-Informed Neural Networks (PINNs) [24, 59, 63], the Deep Ritz method [25], and related variants [41, 7, 32, 8, 22]. These methods hold the promise of providing mesh-free approximations of PDE solutions and of exploiting the dimension-independent approximation power of neural networks [16]. However, their practical adoption is often hindered by optimization difficulties. First-order methods like Adam [42] typically perform poorly, plateauing at low accuracy [46, 67, 60, 68]. This limitation is generally attributed not to the expressiveness of neural networks, but to the ill-conditioning of the loss [60, 21, 46]. Consequently, specialized optimization strategies are increasingly being advocated [19, 55, 62].
@@
More principled second-order optimization algorithms have only recently begun to emerge [54, 55, 21, 73, 20, 60, 33, 38]. Among them, (Energy) Natural Gradient Descent (NGD) [54] has shown remarkable performance, achieving high accuracy in few iterations by using a search direction of the form $\mathbf{d}=-(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})^{-1}\nabla L(\boldsymbol{\theta})$, where $\boldsymbol{\theta}\in\mathbb{R}^{p}$ denotes the neural network parameters, $L$ is the loss function, $\mathbf{G}(\boldsymbol{\theta})$ the positive semidefinite Gramian matrix (see Section 2.2), and $\mu>0$ a regularization parameter. However, solving linear systems with $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ using direct methods has a cubic cost $\mathcal{O}(p^{3})$, making it impractical for large-scale networks. To mitigate this, matrix-free strategies have been proposed [38, 73], where the system is solved approximately using the conjugate gradient (CG) method using only matrix-vector products (matvecs) with $\mathbf{G}(\boldsymbol{\theta})$, without assembling the matrix. In practice, however, $\mathbf{G}(\boldsymbol{\theta})$ is often ill-conditioned, as shown in Figure 1, leading to slow CG convergence and motivating the need for preconditioning. In this work, we address this issue using tools from Randomized Numerical Linear Algebra (RandNLA) [51, 53]. Our main contributions can be summarized as follows:
  • We demonstrate how to efficiently compute matrix-vector products with the Gramian in a more general setting than those considered in [55, 38, 73].
  • We propose the use of randomized Nyström preconditioning [26] to construct a preconditioner for $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ via a low-rank approximation of the Gramian. This turns the strong spectral decay of the Gramian from a challenge into an asset for our approach.
  • We introduce a novel algorithm, NyströmNGD, that combines matrix-free NGD with Randomized Nyström preconditioning.
  • We present results showing significant improvements over prior NGD-based methods.
Related works
@@
Various strategies have been proposed to address training pathologies in PINNs and other neural network-based PDE solvers. These include adaptive quadrature point selection and loss term weighting [67, 13, 66, 69, 71, 68]. More recently, second-order optimizers derived from preconditioning the underlying infinite-dimensional PDE operator have shown promise [21, 55, 54, 33]. Natural Gradient Descent (NGD) has emerged as a unifying framework for these methods, as each can be seen as NGD under a different choice of the optimization metric [56, 55]. Matrix-free implementations have been explored for specific metrics [38, 73], but none of these works addresses the ill-conditioning of the Gramian, which hampers CG convergence in practice. To address this bottleneck, our work bridges NGD with Randomized Numerical Linear Algebra (RandNLA). The use of RandNLA in Machine Learning is well-established [27, 23, 65]. Randomized Nyström preconditioning [26] has been applied to PINNs in [60], but only as a final stage after Adam and L-BFGS, and applying the Nyström approximation, designed for symmetric positive semidefinite matrices, to the Hessian, which may be indefinite. Alternative notions of preconditioning, where the preconditioner is embedded into the loss function, are explored in [47, 7]. K-FAC has also been used to approximate the Gramian [20], but not as a preconditioner, and it requires problem-specific adaptations.
@@
Figure 1: Training of a PINN for the 3D Poisson problem with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$, using Energy Natural Gradient Descent [54]. Left: spectral decay of the Gramian $\mathbf{G}(\boldsymbol{\theta})$ at different iterations. Right: $\mathrm{H}^{1}(\Omega)$ relative error during training.
@@ 2.2 Natural Gradient Descent
Consider the minimization problem on a Riemannian manifold $(\mathcal{M},g)$
$\inf_{u\in\mathcal{M}}E(u).$ (1)
To approximate its solution, we introduce a parametrized ansatz $\mathcal{N}=\left\{u_{\boldsymbol{\theta}}\coloneqq\rho(\boldsymbol{\theta})\,:\,\boldsymbol{\theta}\in\Theta\right\}\subseteq\mathcal{M}$, where $\rho:\Theta\to\mathcal{M}$ is a smooth parametrization map. We then consider the optimization problem
$\inf_{\boldsymbol{\theta}\in\Theta}L(\boldsymbol{\theta})\coloneqq\inf_{\boldsymbol{\theta}\in\Theta}E(u_{\boldsymbol{\theta}})=\inf_{\boldsymbol{\theta}\in\Theta}(E\circ\rho)(\boldsymbol{\theta}).$
For instance, $u_{\boldsymbol{\theta}}$ could be a neural network with parameters $\boldsymbol{\theta}$, though the framework applies more generally. We will refer to $\Theta\subseteq\mathbb{R}^{p}$ as the parameter space, and to $\mathcal{M}$ as the function space.
Natural Gradient Descent (NGD) [3, 50, 56, 55] follows the steepest descent direction according to the metric on $\mathcal{M}$, the “natural” one for the function $u_{\boldsymbol{\theta}}$, and pulls it back to the parameter space. Let $\boldsymbol{\theta}$ be the current iterate. The steepest descent direction in function space would be $-\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})$. In NGD, the corresponding the parameter update direction is $-\mathbf{d}$, where the natural gradient $\mathbf{d}\in\mathbb{R}^{p}$ is such that the first-order expansion $u_{\boldsymbol{\theta}-\alpha\mathbf{d}}\approx u_{\boldsymbol{\theta}}-\alpha\mathrm{D}\rho(\boldsymbol{\theta})[\mathbf{d}]$ best approximates (in the metric $g$) the update in function space. This corresponds to $\mathrm{D}\rho(\boldsymbol{\theta})[\mathbf{d}]=\sum_{j=1}^{p}d_{j}\cdot\partial_{\theta_{j}}u_{\boldsymbol{\theta}}$ being the $g_{u_{\boldsymbol{\theta}}}$-orthogonal projection of $\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})$ onto the model’s generalized tangent space $\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}=\mathrm{Span}\left(\left\{\partial_{\theta_{1}}u_{\boldsymbol{\theta}},\dots,\partial_{\theta_{p}}u_{\boldsymbol{\theta}}\right\}\right)\subset\mathrm{T}_{u_{\boldsymbol{\theta}}}\mathcal{M}.$ As shown, e.g, in [3, 56], the natural gradient is given by $\mathbf{d}=\mathbf{G}(\boldsymbol{\theta})^{\dagger}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta})$, where $\mathbf{G}(\boldsymbol{\theta})^{\dagger}$ is the Moore–Penrose pseudoinverse of the matrix
$\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}={g_{u_{\boldsymbol{\theta}}}}\left({\partial_{\theta_{i}}u_{\boldsymbol{\theta}}},\,{\partial_{\theta_{j}}u_{\boldsymbol{\theta}}}\right),$ (2)
often referred to as the Gramian [72, 55] or Fisher information matrix [3, 50]. As observed in the literature [72, 12], it is often beneficial to add a damping term to improve numerical stability and convergence, yielding the regularized direction

### v3

# Accelerating Natural Gradient Descent for PINNs with Randomized Nyström Preconditioning

@@ Abstract
Natural Gradient Descent (NGD) has emerged as a promising optimization algorithm for training neural network-based solvers for partial differential equations (PDEs), such as Physics-Informed Neural Networks (PINNs). However, its practical use is often limited by the high computational cost of solving linear systems involving the Gramian matrix. While matrix-free NGD methods based on the conjugate gradient (CG) method avoid explicit matrix inversion, the ill-conditioning of the Gramian significantly slows the convergence of CG. In this work, we extend matrix-free NGD to broader classes of problems than previously considered and propose the use of Randomized Nyström preconditioning to accelerate convergence of the inner CG solver. The resulting algorithm demonstrates substantial performance improvements over existing NGD-based methods and other state-of-the-art optimizers on a range of PDE problems discretized using neural networks.
Key words: Natural Gradient, Nyström Preconditioning, Neural Networks, PDEs
@@ 1 Introduction
Neural network (NN) based approaches for solving partial differential equations (PDEs) have recently attracted significant attention as promising alternatives to traditional numerical methods for both forward and inverse problems. A common strategy is to incorporate the governing PDE directly into the loss function, as in Physics-Informed Neural Networks (PINNs) [29, 70, 75], Variational PINNs [49, 50], the Deep Ritz method [30], and various subsequent extensions [8, 38, 9, 25, 73]. These methods hold the promise of providing mesh-free approximations of PDE solutions and of exploiting the approximation power of neural networks [28, 23], particularly in high dimensions [18], positioning them as promising tools for high-dimensional, inverse, or complex domain problems. Despite this theoretical promise, the practical adoption and reliability of NN-based PDE solvers are often hindered by significant optimization challenges. Standard first-order algorithms, notably Adam [51], have been reported to often perform poorly, stagnating at solutions with low accuracy [55, 80, 71, 81]. This issue is generally attributed not to the expressiveness of neural networks, but to the ill-conditioning of the loss landscape [71, 24, 55]. Consequently, specialized optimization strategies tailored to PDE-related loss functions are increasingly being advocated [21, 65, 74].
To counteract these training pathologies, various strategies have been proposed. A body of work focuses on loss weighting schemes or adaptive selection of collocation points to balance contributions from different terms in the loss and reduce quadrature errors [80, 14, 79, 82, 84]. While these techniques can mitigate some failure modes, they do not fully resolve the underlying ill-conditioning introduced by discretizing a PDE operator. Only recently, more principled second-order optimization algorithms have begun to emerge [64, 65, 24, 86, 22, 71, 39, 46], often derived from preconditioning the underlying infinite-dimensional PDE operator [24, 65, 64, 39]. Among them, Natural Gradient Descent (NGD) has emerged as a particularly promising and unifying framework, as several earlier approaches can be interpreted as NGD with different choices of the optimization metric [66, 65].
@@
NGD has demonstrated remarkable performance, achieving high accuracy in few iterations by using a search direction of the form $\mathbf{d}=-(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})^{-1}\nabla L(\boldsymbol{\theta})$, where $\boldsymbol{\theta}\in\mathbb{R}^{p}$ denotes the neural network parameters, $L$ is the loss function, $\mathbf{G}(\boldsymbol{\theta})$ the positive semidefinite Gramian matrix (see Section˜3.1), and $\mu>0$ a regularization parameter. However, solving linear systems with $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ via direct methods incurs a cubic cost $\mathcal{O}(p^{3})$ in the number of network parameters, making NGD impractical for large-scale networks. To alleviate this, matrix-free strategies have been proposed [46, 86], where the system is solved approximately using the conjugate gradient (CG) method using only matrix-vector products (matvecs) with $\mathbf{G}(\boldsymbol{\theta})$, without assembling the matrix. In practice, however, $\mathbf{G}(\boldsymbol{\theta})$ is often ill-conditioned, as shown in Figure˜1, leading to slow CG convergence and motivating the need for preconditioning. In this work, we address this issue using tools from Randomized Numerical Linear Algebra (RandNLA) [60, 63].
The core contributions of this work are as follows. First, we provide a general methodology to efficiently compute matrix-vector products with the Gramian using automatic differentiation, extending beyond the specific metrics considered in prior works [65, 46, 86]. Second, we propose the use of randomized Nyström preconditioning [31] to construct a preconditioner for $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$. Derived from a low-rank approximation of the Gramian, this preconditioner leverages the empirically observed strong spectral decay, turning it from a challenge into an asset. Third, we introduce a novel optimization algorithm for neural network-based PDE solvers, termed NyströmNGD, which combines matrix-free NGD with Randomized Nyström preconditioning. Finally, we present extensive numerical experiments demonstrating significant acceleration and improved final accuracy compared to existing NGD-based methods and other state-of-the-art optimizers.
@@
Matrix-free NGD implementations have already been explored for specific metrics in [46, 86], but none of these works address the ill-conditioning of the Gramian, which hampers CG convergence in practice. To address this bottleneck, our work bridges NGD with Randomized Numerical Linear Algebra (RandNLA), whose use in Machine Learning is well-established [32, 27, 78]. Randomized Nyström preconditioning [31] has been applied to PINNs in [71], but only as a final stage after Adam and L-BFGS, and applying the Nyström approximation, designed for symmetric positive semidefinite matrices, to the Hessian, which may be indefinite. RandNLA has been applied to NGD in the contemporaneous and independent work [61], but there RandNLA is used in the so-called sketch-and-solve paradigm, whereas we employ it to construct a preconditioner. Moreover, in that work the Nyström approximation is not employed even if the Gramian is symmetric positive semidefinite, and the approach is not fully matrix-free. Alternative notions of preconditioning, where the preconditioner is embedded into the loss function, are explored in [56, 8]. Kronecker-Factored Approximate Curvature (K-FAC) [22] has also been used to approximate the Gramian [22], but not as a preconditioner, and it requires problem-specific adaptations.
Paper outline
The remainder of the paper is structured as follows. In Section˜2, we introduce the abstract formulation of the problem and the neural network framework. Section˜3 reviews Natural Gradient Descent and presents its matrix-free formulation. In Section˜4, we provide the necessary background on the randomized Nyström approximation, and in Section˜4.2 we present our proposed algorithm, NyströmNGD, in full detail. Section˜5 contains numerical experiments on a range of PDE-related problems, comparing our method against state-of-the-art optimizers. Finally, Section˜6 summarizes our findings and outlines directions for future research.
@@
Figure 1: Spectral decay of the Gramian matrix when training a PINN for the 3D Poisson problem with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$, using Energy Natural Gradient Descent [64]. Left: spectral decay of the Gramian $\mathbf{G}(\boldsymbol{\theta})$ at different iterations, with each color corresponding to a different iteration. Right: $\mathrm{H}^{1}(\Omega)$ relative error during training.
@@ 2.1 Notation
Hilbert and Banach spaces are indicated using roman capital letters. Linear operators are denoted with calligraphic letters. Vectors in $\mathbb{R}^{n}$ are in boldface, and matrices use boldface capital letters.
@@
2.2 Abstract Setup
We consider a general minimization problem of the form
$\inf_{u\in\mathrm{V}}E(u),$ (1)
where $\mathrm{V}$ is a Hilbert space endowed with a possibly point-dependent metric $g$ which gives to $\mathrm{V}$ the structure of a Riemannian manifold, and $E:\mathrm{V}\to\mathbb{R}$ is a smooth function, referred to as the energy.
To approximate its solution, we introduce a parametrized ansatz
$\mathcal{N}=\left\{u_{\boldsymbol{\theta}}\coloneqq\rho(\boldsymbol{\theta})\,:\,\boldsymbol{\theta}\in\Theta\right\}\subseteq\mathrm{V},$
where $\rho:\Theta\to\mathrm{V}$ is a smooth map and $\Theta\subseteq\mathbb{R}^{p}$ is open. We then consider the finite-dimensional optimization problem
$\inf_{\boldsymbol{\theta}\in\Theta}L(\boldsymbol{\theta})\coloneqq\inf_{\boldsymbol{\theta}\in\Theta}E(u_{\boldsymbol{\theta}})=\inf_{\boldsymbol{\theta}\in\Theta}(E\circ\rho)(\boldsymbol{\theta}).$ (2)
Here, $u_{\boldsymbol{\theta}}$ could be a neural network with parameters $\boldsymbol{\theta}$, though the framework applies more generally. We will refer to $\Theta\subseteq\mathbb{R}^{p}$ as the parameter space, and to $\mathrm{V}$ as the function space. Note that while the original optimization problem (1) may be infinite dimensional, the ansatz is parametrized by a finite-dimensional $\boldsymbol{\theta}\in\mathbb{R}^{p}$.
2.2.1 Discretization of PDEs
The focus of this article is on problems arising from the discretization of PDEs. In this case, $\mathrm{V}$ is a Hilbert space of functions defined on a domain $\Omega\subset\mathbb{R}^{d}$, e.g., $\mathrm{V}=\mathrm{H}^{1}(\Omega)$, and $E$ is derived from the strong or weak formulation of a PDE on $\Omega$. This abstract setting (2) encompasses, for instance, the Deep Ritz method [30], PINNs [70, 29], Variational PINNs (VPINNs)[49], and other PINNs variants [50, 8, 38, 25, 73].
As a general framework for PDEs we consider the following. Let $\mathrm{V}$ be Hilbert space and let $\mathrm{X},\mathrm{Y}$ be Banach spaces. We seek $u\in\mathrm{V}$ such that
$\begin{cases}\mathcal{F}(u)=f,&\text{in }\Omega,\\ \mathcal{B}(u)=g,&\text{on }\partial\Omega,\end{cases}$ (3)
where $\mathcal{F}:\mathrm{V}\to\mathrm{X}$ is a differential operator, $f\in\mathrm{X}$ is the source term, $\mathcal{B}:\mathrm{V}\to\mathrm{Y}$ is a boundary operator, and $g\in\mathrm{Y}$ is the boundary data. The energy can then be defined as
$E(u)=\frac{1}{2}\left\|{\mathcal{F}(u)-f}\right\|_{\mathrm{X}}^{2}+\frac{1}{2}\left\|{\mathcal{B}(u)-g}\right\|_{\mathrm{Y}}^{2},$
where possibly weighting factors can be included, and data misfit terms can be added.
The choice of function spaces and well-posedness depend on the specific PDE and on whether a strong or weak formulation is used. For example, consider the Poisson equation with homogeneous Dirichlet boundary conditions:
$\begin{cases}-\Delta u=f,&\text{in }\Omega,\\ u=g,&\text{on }\partial\Omega.\end{cases}$ (4)
For a strong formulation, once can employ the spaces $\mathrm{V}=\mathrm{H}^{2}(\Omega)$, $\mathrm{X}=\mathrm{L}^{2}(\Omega)$, and $\mathrm{Y}=\mathrm{L}^{2}(\partial\Omega)$, with energy
$E(u)=\frac{1}{2}\left\|{-\Delta u-f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2}.$
If instead a weak formulation is used, the natural spaces are $\mathrm{V}=\mathrm{H}^{1}_{0}(\Omega)$, $\mathrm{X}=\mathrm{V}^{\prime}=\mathrm{H}^{-1}(\Omega)$, and $\mathrm{Y}=\mathrm{H}^{1/2}(\partial\Omega)$. Given a lifting $\tilde{g}\in\mathrm{H}^{1}(\Omega)$ of the Dirichlet data, the energy is
$E(u)=\frac{1}{2}\left\|{\mathcal{A}u-\tilde{f}}\right\|_{\mathrm{H}^{-1}(\Omega)}^{2},\qquad\text{where}\qquad\left\langle{\mathcal{A}u},\,{v}\right\rangle_{\mathrm{H}^{-1}\times\mathrm{H}^{1}_{0}}=\int_{\Omega}\nabla u\cdot\nabla v,\quad\tilde{f}=f-\mathcal{A}\tilde{g}$.
2.3 Neural Networks
Neural networks (NNs) serve as the primary parametrized ansatz in this work. For simplicity, we focus on feedforward networks, though our methods are architecture-agnostic and apply to other network types as well. A neural network with input dimension $d$ and output dimension $d^{\prime}$ is a function $u_{\boldsymbol{\theta}}^{\mathcal{NN}}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ defined as the composition affine and nonlinear transformations. Let $L$ be the number of hidden layers, let $n_{0}=d,n_{L+1}=d^{\prime}$, and $n_{1},\dots,n_{L}$ be the number of neurons in each hidden layer, and let $\sigma$ be a nonlinear activation function. The NN is defined as:
$\begin{split}\mathbf{x}_{0}&=\mathbf{x}\in\mathbb{R}^{d},\\ \mathbf{x}_{\ell}&=\sigma(\mathbf{W}_{\ell}\mathbf{x}_{\ell-1}+\mathbf{b}_{\ell})\in\mathbb{R}^{n_{\ell}},\hskip 18.49988pt\text{for }\ell=1,\dots,L,\\ u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x})=\mathbf{x}_{L}&=\mathbf{W}_{L+1}\mathbf{x}_{L}+\mathbf{b}_{L+1}\in\mathbb{R}^{d^{\prime}},\end{split}$
where $\mathbf{W}_{\ell}\in\mathbb{R}^{n_{\ell}\times n_{\ell-1}}$ and $\mathbf{b}_{\ell}\in\mathbb{R}^{n_{\ell}}$ are the weights and biases of layer $\ell$, respectively. The parameters of the network are $\boldsymbol{\theta}=[\mathbf{W}_{1},\mathbf{b}_{1},\dots,\mathbf{W}_{L+1},\mathbf{b}_{L+1}]\in\mathbb{R}^{p}$, stacked into a single vector, with $p=\sum_{\ell=1}^{L+1}(n_{\ell}n_{\ell-1}+n_{\ell})$. No activation is applied in the last layer.
2.3.1 Finite Element Interpolated Neural Networks (FEINNs)
While standard NNs are typically used with collocation-based methods like PINNs, to deal with variational formulations of PDEs, we consider Finite Element Interpolated Neural Networks (FEINNs) [8, 7, 10], where a neural network is interpolated onto a finite element (FE) space.
Let $\Omega\subset\mathbb{R}^{d}$ be a Lipschitz polyhedral domain with a conforming simplicial mesh $\mathcal{T}_{h}$, and let $\mathrm{V}_{h}\subset\mathrm{V}=\mathrm{H}^{1}(\Omega)$ be the FE space with $\mathrm{V}_{h,0}=\mathrm{V}_{h}\cap\mathrm{H}^{1}_{0}(\Omega)$. Denote the Lagrangian basis by $\{\varphi_{i}\}_{i=1}^{N}$, where $\varphi_{i}$ corresponds to node $\mathbf{x}_{i}$, and let $\Pi_{h}:C^{0}(\Omega)\to\mathrm{V}_{h,0}$ be the FE interpolation operator obtained by evaluation at interior nodes. Given a discrete lifting $\bar{u}_{h}$ of Dirichlet data, the FEINN is defined as
$u_{\boldsymbol{\theta}}=\bar{u}_{h}+\Pi_{h}u_{\boldsymbol{\theta}}^{\mathcal{NN}}=\bar{u}_{h}n+\sum_{i\,:\,\mathbf{x}_{i}\in\mathring{\Omega}}u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x}_{i})\varphi_{i}.$
For simplicity, we assumed that Dirichlet boundary conditions are imposed on the whole boundary, but mixed boundary conditions can also be accommodated in the framework. Further details are provided in [8, 7].
This approach offers several advantages. First, Dirichlet boundary conditions can be approximately enforced by applying them to the FE interpolant using standard FE techniques, avoiding the need to impose them directly on the neural network output [10, 11]. Second, it alleviates integration challenges, a known issue in neural network-based PDE solvers [72], especially in the Deep Ritz method [76]. Finally, spatial derivatives can be computed at the FE level, eliminating the need for automatic differentiation with respect to the input of the neural network.
3 Matrix-free Natural Gradient Descent
@@ 3.1 Natural Gradient Descent
The Natural Gradient Descent (NGD) method [3, 59, 66, 65] targets optimization problems of the form (2) following the first-optimize-then-discretize paradigm. NGD moves along the steepest descent direction according to the metric on $\mathrm{V}$, the “natural” one for the function $u_{\boldsymbol{\theta}}$, and pulls it back to the parameter space.
Let $\boldsymbol{\theta}$ be the current iterate. The steepest descent direction in function space is $d=-\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})$, the Riemannian gradient of $E$ in the metric $g$, and the corresponding update
$u\leftarrow u_{\boldsymbol{\theta}}-\alpha\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}}).$
In NGD, the parameter update direction is $\mathbf{d}$, where the natural gradient $\mathbf{d}\in\mathbb{R}^{p}$ is such that the first-order expansion
$u_{\boldsymbol{\theta}-\alpha\mathbf{d}}\approx u_{\boldsymbol{\theta}}-\alpha\mathrm{D}\rho(\boldsymbol{\theta})[\mathbf{d}]=u_{\boldsymbol{\theta}}-\alpha\sum_{j=1}^{p}d_{j}\cdot\partial_{\theta_{j}}u_{\boldsymbol{\theta}},$
best approximates (in the metric $g$) the update in function space, i.e.,
$\mathbf{d}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathbb{R}^{p}}\left\|{\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})-\mathrm{D}\rho(\boldsymbol{\theta})[\mathbf{w}]}\right\|_{g_{u_{\boldsymbol{\theta}}}}.$
This corresponds to the $g_{u_{\boldsymbol{\theta}}}$-orthogonal projection of $\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})$ onto the model’s generalized tangent space
$\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}=\mathrm{Span}\left(\left\{\partial_{\theta_{1}}u_{\boldsymbol{\theta}},\dots,\partial_{\theta_{p}}u_{\boldsymbol{\theta}}\right\}\right)\subset\mathrm{V}.$
If $\partial_{\theta_{1}}u_{\boldsymbol{\theta}},\dots,\partial_{\theta_{p}}u_{\boldsymbol{\theta}}$ are not linearly independent, the minimum is not unique, and $\mathbf{d}$ of minimal norm is taken. As shown, e.g, in [3, 66], the natural gradient can be computed as
$\mathbf{d}=\mathbf{G}(\boldsymbol{\theta})^{\dagger}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta}),\qquad\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}={g_{u_{\boldsymbol{\theta}}}}\left({\partial_{\theta_{i}}u_{\boldsymbol{\theta}}},\,{\partial_{\theta_{j}}u_{\boldsymbol{\theta}}}\right),$ (5)
where $\mathbf{G}(\boldsymbol{\theta})\in\mathbb{R}^{p\times p}$ is often referred to as the Gramian [85, 65] or Fisher information matrix [3, 59], and $\mathbf{G}(\boldsymbol{\theta})^{\dagger}$ denotes its Moore–Penrose pseudoinverse. As observed in the literature [85, 13], it is often beneficial to add a damping term to improve numerical stability and convergence, yielding the regularized direction